OpenStudy (anonymous):

solve for x and y 2x-y=2 and y=(x-2)(x-1)

OpenStudy (diyadiya):

\[y=(x-2)(x-1)= x^2-x-2x+2 =x^2 -3x+2 \] Plug in the equation of y in 2x-y=2 \[2x-(x^2 -3x+2) =2\]\[2x-x^2+3x-2=2\]\[-x^2+5x-4=0\]

OpenStudy (diyadiya):

Can you solve this quadratic equation ?

OpenStudy (anonymous):

i tried using the formula to get the value of x but i get math error

OpenStudy (diyadiya):

You can do it without using a calculator :D

OpenStudy (anonymous):

im not very good at factoring thats why i use the formula

OpenStudy (diyadiya):

Factorise -x^2+5x -4 =0 For easier calculation multiply the whole equation by -1 \(x^2-5x+4=0 \) You can use the quadratic formula to solve it \[\huge x= \frac{-b \pm \sqrt{b^2 -4ac}}{2a}\]

OpenStudy (diyadiya):

Here ; a = 1 b = -5 c= 4

OpenStudy (diyadiya):

Plug in the values in the formula and find x When you've found out x, put the value of x in any of the equation in the question & find y tell me if you didn't understand any of the above steps

OpenStudy (anonymous):

x=6.97 and -1.97 is that corret for the value of x

OpenStudy (diyadiya):

No! Can you show me how you did that?

OpenStudy (anonymous):

i used the formula